makaylab:

y=-x+1 y=-1/4x-2 solve

makaylab:

plz help

makaylab:

1 attachment
minustempo:

Since y = -x+1, you can plug that into the second equation -x+1 = -1/4x - 2 now can you solve it?

makaylab:

@minustempo wrote:
Since y = -x+1, you can plug that into the second equation -x+1 = -1/4x - 2 now can you solve it?
i'll try

makaylab:

@minustempo wrote:
Since y = -x+1, you can plug that into the second equation -x+1 = -1/4x - 2 now can you solve it?
nope

surjithayer:

\[y=x-1\] \[y=-\frac{ 1 }{ 4 }x-2\] substitute the value of y in the second equation \[x-1=-\frac{ 1 }{ 4 }x-2\] multiply by 4 4x-4=-x-8 4x+x=-8+4 5x=-4 x=?

surjithayer:

then put the value of x in either eqution and find the value of y.

LostLegacy06:

wait are you solving or graphing? like do you just have to graph it

makaylab:

1 attachment
LostLegacy06:

so you just have to graph it I dont think you have to solve them they are already in y=mx+b form

makaylab:

i need help tho...

LostLegacy06:

alr alr one sec

makaylab:

ok

LostLegacy06:

ok so we're going to start with y=-1/4+2

LostLegacy06:

just like put a dot at y=2 since the y intercept is 2

LostLegacy06:

now you go down 1 and over 4 to the right

makaylab:

k. done that and have to solve now

LostLegacy06:

1 attachment
LostLegacy06:

what exactly are you solving they are both in the correct form

makaylab:

the 2 is -

makaylab:

wait now i confused

LostLegacy06:

so am i lol the 2 is negative?

LostLegacy06:

oh snap one sec

makaylab:

1 attachment
LostLegacy06:

1 attachment
LostLegacy06:

ok it should look like that

LostLegacy06:

now we have to graph y = -x + 1

makaylab:

ok

LostLegacy06:

1 attachment
makaylab:

kk